# Celestial Topology, Symmetry Theories, and Evidence for a Non-SUSY D3-Brane CFT

### v1

@@ 2 Causal Structures in a Symmetry Theory
Along these lines, suppose we have a string background which engineers a QFTD on $D$-dimensional Minkowski space, as obtained from a background of the form $\mathbb{R}^{D-1,1}\times X$ with metric of the form: [^11] Of course, in actual compactifications it is sometimes necessary to multiply the $\mathbb{R}^{D-1,1}$ and $X$ directions of the metric by suitable warp factors, which due to the various instabilities will also be accompanied by both $r$ and $t$ dependence. However, to investigate the causal structure of the various instabilities and their impact on the celestial topology, it suffices to make the approximation used below.
@@ 3.1 The $\lambda\ll 1$ Regime
At this point, let us address one particularly subtle aspect of our analysis so far; we are not simply considering an excited state of the QFT and watching its trajectory as a function of time. In that case we would observe an oscillation where the fields move to the minimum of the effective potential and then move back up. Rather, what is happening is that as we roll from the initial excited state down to the minimum of the effective potential, some energy is being radiated away as closed string modes; this is simply the tachyonic shell moving away from the D3-brane. Consequently, if we continue to work purely about the ground state specified by the $\mathbb{R}^{6}/\Gamma^{\prime}$ geometry we will instead perceive a “vertical wall” in the associated SymTh. This is the point of view adopted in [39]; here we are focusing on how the closed string dynamics is encoded in the D3-brane worldvolume theory.
@@
Observe that as we increase $\lambda$ (but still keeping $\lambda\ll 1$), the magnitude of $M^{2}$ decreases. Thus, the initial speed of the transmission in the SymTh is slower than the speed of light.161616“Tachyonic” simply means we have an unstable configuration when $M^{2}<0$. Nothing is moving faster than the speed of light. Increasing the magnitude of $M^{2}$ in this regime just makes the effective potential steeper and steeper. Conversely, decreasing the magnitude of $M^{2}$ makes the effective potential more shallow. On the other hand, at late times, we know that the instability is carried out by a massless metric and dilaton pulse, and so asymptotically approaches the speed of light. Putting these facts together, we conclude that in the Penrose diagram, the eventual collision of the trajectory in the $(T,R)$ directions with $\mathcal{I}^{+}$ continues to move closer to $i^{+}$ as we increase $\lambda$ (see figure 4). In the interpolating region between $S^{5}/\Gamma$ and $S^{5}/\Gamma^{\prime}$ there is a bordism which connects the two pieces of $\mathcal{I}^{+}$.
@@ 3.2 The $\lambda\gg 1$ Regime
Nevertheless, taking the near horizon geometry at face value, we can proceed as in [44]. The appearance of an $S^{5}/\Gamma$ factor means that in the “branes at infinity” picture we have the corresponding topological symmetry operators induced directly from the topology of this geometry. Moreover, the original tachyonic mode of the twisted sector has been stretched to a large value, and is no longer present [57, 35, 44]. On the other hand, the appearance of a non-perturbative bubble destroys the solution at finite proper time; the putative CFT dual seems to have disappeared altogether!
@@ 3.3 Symmetry Restoration / Evidence for a Non-SUSY CFT
What happens in the interpolating range of values for $\lambda$? Letting $\ell$ denote a local coordinate on the null ray extending from $i^{0}$ (i.e., $\ell=0$) to $i^{+}$ (i.e., $\ell=1$), we denote by $\ell(\lambda)$ the value of $\ell$ at which the instability of the SymTh intersects $\mathcal{I}^{+}$. A direct calculation of $\ell(\lambda)$ appears out of reach, but we can nevertheless constrain it based on the topology of the celestial geometry. Note first that in the regime $\lambda\ll 1$, the function $\ell(\lambda)$ is monotonically increasing. On the other hand, it somehow must “turn around” since eventually near $\lambda\gg 1$ the bubble-of-nothing type instability quickly eats the whole spacetime. There is some critical value of $\ell_{\mathrm{closed}}$ beyond which we no longer have an $S^{5}/\Gamma\rightarrow S^{5}/\Gamma^{\prime}$ transition at the boundary, i.e., we cease to have the original closed string tachyon. Likewise, there is some critical value of $\ell_{\mathrm{bubble}}$ which signals the first onset of the bubble-of-nothing type instability, i.e., we instead have a transition $S^{5}/\Gamma\rightarrow\emptyset$.
@@
The considerations here have focused on topological properties of the celestial geometry. It would be interesting to connect these considerations more directly to celestial holography. [^19] See e.g., [62] for a recent review.

### v2

@@ 2 Causal Structures in a Symmetry Theory
Along these lines, suppose we have a string background which engineers a QFTD on $D$-dimensional Minkowski space, as obtained from a background of the form $\mathbb{R}^{D-1,1}\times X$ with metric of the form: [^11] Of course, in actual compactifications it is sometimes necessary to multiply the $\mathbb{R}^{D-1,1}$ and $X$ directions of the metric by suitable overall warp factors, which due to the various instabilities will also be accompanied by both $r$ and $t$ dependence. However, to investigate the causal structure of the various instabilities and their impact on the celestial topology, it suffices to make the approximation used below. We emphasize that in the present context, the warp factors under considerations are those associated with probe D3-branes. As such, the asymptotic causal structure of the spacetime geometry is unaffected. This is important because we will shortly use the boundary topology to analyze symmetry breaking patterns, much as has already been carried out quite successfully in the supersymmetric setting.
@@ 3.1 The $\lambda\ll 1$ Regime
At this point, let us address one particularly subtle aspect of our analysis so far; we are not simply considering an excited state of the QFT and watching its trajectory as a function of time. In that case we would observe an oscillation where the fields move to the minimum of the effective potential and then move back up. Rather, what is happening is that as we roll from the initial excited state down to the minimum of the effective potential, some energy is being radiated away as closed string modes; this is simply the tachyonic shell moving away from the D3-brane. Consequently, if we continue to work purely about the ground state specified by the $\mathbb{R}^{6}/\Gamma^{\prime}$ geometry we will instead perceive a “vertical wall” in the associated SymTh. This is the point of view adopted in [39]; here we are focusing on how the closed string dynamics is encoded in the D3-brane worldvolume theory. Again, the point here is that we are taking at face value the UV non-supersymmetric quiver gauge theory description and simply tracking what happens at long distance scales.
@@
Observe that as we increase $\lambda$ (but still keeping $\lambda\ll 1$), the magnitude of $M^{2}$ decreases. Thus, the initial speed of the transmission in the SymTh is slower than the speed of light.161616“Tachyonic” simply means we have an unstable configuration when $M^{2}<0$. Nothing is moving faster than the speed of light. Increasing the magnitude of $M^{2}$ in this regime just makes the effective potential steeper and steeper. Conversely, decreasing the magnitude of $M^{2}$ makes the effective potential more shallow. On the other hand, at late times, we know that the instability is carried out by a massless metric and dilaton pulse, and so asymptotically approaches the speed of light. Putting these facts together, we conclude that in the Penrose diagram, the eventual collision of the trajectory in the $(T,R)$ directions with $\mathcal{I}^{+}$ continues to move closer to $i^{+}$ as we increase $\lambda$ (see figure 4). In the interpolating region between $S^{5}/\Gamma$ and $S^{5}/\Gamma^{\prime}$ there is a bordism which connects the two pieces of $\mathcal{I}^{+}$. [^17] It is natural to ask whether having $\lambda\neq 0$ simply leads to a runaway potential, i.e., the ground state is simply destroyed altogether. This is incompatible with the closed string picture. Indeed, from [35, 36], we know that the probe D3-branes essentially stay put and, after the tachyonic shell has smoothed away the non-supersymmetric parts of the singularity, eventually the system settles in a supersymmetric geometry probed by the D3-branes. In field theory terms, this means that the system eventually settles in a supersymmetric ground state. We emphasize that this is the behavior in the regime $\lambda\ll 1$.
@@ 3.2 The $\lambda\gg 1$ Regime
Nevertheless, taking the near horizon geometry at face value, we can proceed as in [44]. Ultimately, we will be interested in the asymptotically flat setup, however it is useful to consider this putative $AdS$ dual to characterize the instabilities that might occur away from such a near horizon limit. In this near horizon geometry the appearance of an $S^{5}/\Gamma$ factor means that in the “branes at infinity” picture we have the corresponding topological symmetry operators induced directly from the topology of this geometry. Moreover, the original tachyonic mode of the twisted sector has been stretched to a large value, and is no longer present [57, 35, 44]. On the other hand, the appearance of a non-perturbative bubble destroys the solution at finite proper time; the putative CFT dual seems to have disappeared altogether!
@@ 3.3 Symmetry Restoration / Evidence for a Non-SUSY CFT
What happens in the interpolating range of values for $\lambda$? Our claim is that varying $\lambda$ inevitably leads to a transition which can be pushed all the way to $i^{+}$. To see this, introduce $\ell$ a local coordinate on the null ray extending from $i^{0}$ (i.e., $\ell=0$) to $i^{+}$ (i.e., $\ell=1$), we denote by $\ell(\lambda)$ the value of $\ell$ at which the instability of the SymTh intersects $\mathcal{I}^{+}$. A direct calculation of $\ell(\lambda)$ appears out of reach, but we can nevertheless constrain it based on the topology of the celestial geometry. Note first that in the regime $\lambda\ll 1$, the function $\ell(\lambda)$ is monotonically increasing. On the other hand, it somehow must “turn around” since eventually near $\lambda\gg 1$ the bubble-of-nothing type instability quickly eats the whole spacetime. There is some critical value of $\ell_{\mathrm{closed}}$ beyond which we no longer have an $S^{5}/\Gamma\rightarrow S^{5}/\Gamma^{\prime}$ transition at the boundary, i.e., we cease to have the original closed string tachyon. Likewise, there is some critical value of $\ell_{\mathrm{bubble}}$ which signals the first onset of the bubble-of-nothing type instability, i.e., we instead have a transition $S^{5}/\Gamma\rightarrow\emptyset$.
@@
Let us further emphasize, that indeed in the transition from large to small values of $\lambda$, across $\lambda_{*}$, the potentials displayed in figure 5 are guaranteed to have distinct large field profile, as characteristic for a second order phase transition. For example, a potential with ever rolling solution at small ’t Hooft coupling is excluded by the perturbative analysis at small $\lambda$. There, following [39], we know that after the background relaxes to a stable configuration we are left with a D3-brane probing a supersymmetric geometry, i.e., we are guaranteed to settle into a stable vacuum.
The considerations here have focused on topological properties of the celestial geometry. It would be interesting to connect these considerations more directly to celestial holography, though we leave the exploration of this speculation for future endeavors. [^20] See e.g., [62] for a recent review of celestial holography.